Two number cubes were rolled 150 times. Find the experimental probability that a sum less than 4 appeared 15 times
Answer by stanbon(75887)

Two number cubes were rolled 150 times. Find the experimental probability that a sum less than 4 appeared 15 times
----
# of ways to get sum of 2 or 3:: 1+2 = 3
P(get sum of 2 or 3) = 3/36 = 1/12
-------
P(getting sum < 4 in 150 rolls) = 150C3(1/12)^3*(11/12)^147
= binompdf(150,1/12,3) = 0.00089..
Cheers,
Stan H.
--------